A block of mass M is connected to another block of mass M2 by a massless spring of force constant K the blocks are resting on a smooth horizontal plane the blocks are initially at rest and the spring and stretch then a constant force f is applied on block m to pull it find the force on the block M2

Grade:12

1 Answers

Arun
5 years ago

When 2 Blocks are in Contact -



F-f=m_{1}a

f=m_{2}a

- wherein

a=\frac{F}{m_{1}+m_{2}}

f=\frac{m_{2}F}{m_{1}+m_{2}}

Hope it helps

You can ask as many queries on forum. We will be happy to help you.

Regards

Arun

 

 

: Acceleration \: of\: the\: system\: a= \frac{F}{m+M}

Force \: on \: block\: of \: mass\: m= ma= \frac{mF}{m+M}